OpenStudy (anonymous):

derivative of 2xye^-x^2

OpenStudy (anonymous):



myininaya (myininaya):

you have two variable expression so are you doing partial derivatives?

myininaya (myininaya):

if so are you doing the partial w.r.t. x or y?

myininaya (myininaya):

and in your drawing why is there a dy at the end?

myininaya (myininaya):

are we doing the partial w.r.t to y?

myininaya (myininaya):

like is this what you meant? \[\frac{\partial }{ \partial y}(2xye^{-x^2})\]

OpenStudy (anonymous):

yes respect to y

OpenStudy (anonymous):

yes yes @myininaya

myininaya (myininaya):

ok well that means you treat everything by y like a constant

myininaya (myininaya):

so in other words you know \[\frac{d}{dy}cy=c \frac{d}{dy}=c \frac{dy}{dy}=c(1)=c\]

OpenStudy (anonymous):

im confused with this "treat y as a constant" you mean? take out y from the equation?

myininaya (myininaya):

\[2xe^{-x^2} \frac{\partial }{\partial y}y=?\]

myininaya (myininaya):

no treat everything BUT y like a constant

myininaya (myininaya):

that is the only other variable is x treat x like a constant

myininaya (myininaya):

so you see there for your partial i pulled out the constant multiple

myininaya (myininaya):

all you have to do is find the derivative of y w.r.t to y now

OpenStudy (anonymous):

x's the only variable..

OpenStudy (anonymous):

can i use the method udv+vdu?

myininaya (myininaya):

I don't understand. your expression is in terms of x and y. You have two variables. You are finding the partial w.r.t y so that means x is like a constant. so 2x is like a constant here so e^(-x^2) is also like a constant here and the product of constants is a constant so 2xe^(-x^2) is like a constant here

myininaya (myininaya):

\[\frac{\partial }{ \partial y} (2xye^{-x^2})= 2xe^{-x^2} \frac{\partial }{\partial y} y =\text{ you can finish this }\]

OpenStudy (anonymous):

my answer is 2xe^-x^2??